分子病原微生物論 2 単位 (選択)

Molecular Pathogenic Microbiology  長宗秀明 · 教授/ 環境創生工学専攻生命テクノサイエンスコース生物機能工学講座

友安俊文 · 准教授/ 環境創生工学専攻生命テクノサイエンスコース生物機能工学講座

【授業目的】病原微生物の持つ病原因子の分子的特徴やその応用等について最新 の知見を理解する． 【授業概要】病原微生物による感染症に関係する病原因子についての最新の知見 を紹介して課題を出し，ポートフォーリオ方式で課題提出物について評価を 行う． 【授業形式】講義形式とポートフォーリオ形式の併用 【キーワード】病原微生物，病原因子，感染 【先行科目】 [先行科目] 【関連科目】『微生物分子論』 (0.9) 【履修要件】学部教育における微生物学及びその関連科目を理解していること． 【履修上の注意】 [注意] 【到達目標】 1. 病原微生物の病原因子に関する最新の分子論的情報を理解することができる 2. 微生物病原因子の応用技術を分子論的に理解することができる 【授業計画】 1. 細菌毒素 1:CDC 膜孔形成毒素 2. 細菌毒素 2:他の膜孔形成毒素 3. 細菌毒素 3:神経指向性酵素毒素 4. 細菌毒素 4:消化管指向性酵素毒素 5. 細菌毒素 5:蛋白翻訳系指向性酵素毒素 6. 細菌毒素 6:スーパー抗原 7. 抗生物質耐性遺伝子を持つプラスミド類 8. 菌類毒素 1:小分子性毒素 9. 菌類毒素 2:蛋白毒素 10. 他の細菌性病原因子 1:定着因子 11. 他の細菌性病原因子 2:酵素 12. 他の細菌性病原因子 3:分泌系 13. ウイルス病原因子 1:定着因子 14. ウイルス病原因子 2:酵素 15. ウイルス病原因子 3:転写調節関連因子 【成績評価基準】授業日毎に課すレポート 100%で評価する．【教科書】授業中に紹介する【参考書】授業中に紹介する【授業コンテンツ】 http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216881 【対象学生】開講コース学生のみ履修可能【連絡先】

⇒ 長宗 (化生棟 707, 088-656-7525, nagamune@bio.tokushima-u.ac.jp) ����

(オフィスアワー: 月曜日 16:20-17:50) 【備考】各々 2 時間の予習と復習が各 2 時間の授業の理解と単位取得のために必

要である．

Molecular Pathogenic Microbiology 2 units (selection)

Hideaki Nagamune · Professor / Biological Functions, Biological Science and Technology, Earth and Life Environmental Engineering Toshifumi Tomoyasu · Associate Professor / Biological Functions, Biological Science and Technology, Earth and Life Environmental Engineering

Target⟩ This class aims to understand the latest findings on molecular feature and

application of virulence factors expressed in pathogenic microorganisms. Outline⟩ The latest findings on virulence factors concerned with infectious diseases

by pathogenic microorganisms are introduced in this class. Assignments on the introduced topics are set and marked in portfolio system.

Style⟩ Lecture in combination with Portfolio �������⟩ pathogenic microorganism, virulence factor, infection Fundamental Lecture⟩ [先行科目] Relational Lecture⟩ “Biomolecular Engineering”(0.9) Requirement⟩ Students are required to have a good understanding of undergraduate-

level of microbiology and related subjects Notice⟩ [注意]

Goal⟩

1. To understand the latest molecular information on virulence factors of pathogenic microorganisms

2. To understand the application technology of microbial virulence factors in

molecular level

Schedule⟩

1. Bacterial toxins 1:CDC pore-forming toxins

2. Bacterial toxins 2:other pore-forming toxins

3. Bacterial toxins 3:neuron-directed enzymatic toxins

4. Bacterial toxins 4:gastrointestinal tissue-directed enzymatic toxins

5. Bacterial toxins 5:translational step-directed enzymatic toxins

6. Bacterial toxins 6:superantigens

7. Plasmids carrying microbial-resistant genes

8. fungal toxins 1:small molecular toxins

9. fungal toxins 2:protein toxins

10. Other Bacterial virulence factors 1:adherence factors

11. Other Bacterial virulence factors 2:enzymes

12. Other Bacterial virulence factors 3:secretion systems

13. Viral virulence factors 1:adherence factors

14. Viral virulence factors 2:enzymes 15. Viral virulence factors 3:transcriptional factors Evaluation Criteria⟩ Assignments count 100% Textbook⟩ To be introduced in the class Reference⟩ To be introduced in the class

Contents⟩ http://cms.db.tokushima-u.ac.jp/cgi-bin/toURL?EID=216881

Student⟩ Able to be taken by only specified class(es)

Contact⟩

⇒ Nagamune (G707, +81-88-656-7525, nagamune@bio.tokushima-u.ac.jp)

���� (Office Hour: Monday 16:20-17:50)

Note⟩ Preparation and review of two hours each are required for understanding of every class with two hours and taking the credits.